# Origen: Zero-Shot 3D Orientation Grounding in Text-to-Image Generation

### v1

@@ Abstract
Figure 1: 3D orientaion-grounded text-to-image generation results of Origen. We present Origen, the first zero-shot method for 3D orientation grounding in text-to-image generation across multiple objects and diverse categories. Origen generates high-quality images that are accurately aligned with the grounding orientation conditions, indicated by the colored arrows, and the input text prompts.
**footnotetext: Equal contribution.
@@ 1 Introduction
Since no existing method has quantitatively evaluated 3D orientation grounding in text-to-image generation (except for user studies by Cheng et al. [11]), we curate a benchmark based on the MS-COCO dataset [42], mixing and matching object classes and orientations to create images with single or multiple orientation-grounded objects. We demonstrate that Origen significantly outperforms previous orientation-conditioned image generative models [11, 46] on both our benchmark and user studies. Since prior models cannot condition on multiple objects (whereas Origen can, as shown in Fig. 1), comparisons are conducted under single-object conditioning. Additionally, we perform experiments to further validate the superior performance of our method over text-to-image generative models with orientation-specific prompts and other training-free guided sampling strategies.
@@
  • We also propose a reward-adaptive time rescaling method that further accelerates convergence.
@@ 2 Related Work
A number of training-free methods have been proposed for guided generation tasks. DPS [13], MPGD [20], and PiGDM [58] update the noisy data point at each step using a given reward function. Yu et al. [74] and Ye et al. [72] take this further by introducing rewinding, where intermediate data points are regenerated by reversing the generative denoising process. The core principle of this approach is leveraging the conditional expectation of the clean image from a noisy image at an intermediate step. The expected future reward can also be computed from the expected clean image, allowing gradient ascent to update the noisy image. However, a key limitation of these methods is that the expected clean image from a diffusion model is often too blurry to accurately predict future rewards. While distilled or fine-tuned diffusion [59, 27, 41] and flow models [48, 49] can mitigate this issue, their straightened trajectories lead to insufficiently small updates to the noisy image during gradient ascent at intermediate timesteps.
Recent approaches [5, 64, 19, 16] attempt to address this limitation by updating the initial noise rather than the intermediate noise. Notably, Eyring et al. [16] introduce a one-step generative model to efficiently iterate the initial noise update through one-step generation and future reward computation. However, gradient ascent with respect to the initial noise often suffers from local optima and leads to deviations from real images, even with heuristic regularization [55]. To overcome this limitation, we propose a novel sampling-based approach rather than an optimization-based one, leveraging Langevin dynamics to effectively balance reward maximization and realism while maintaining a simple implementation.
@@
For controlling image generation, ControlNet [75] and IP-Adapter [71] are commonly used to utilize a pretrained generative model to control for various conditions, though they require training data. For 3D orientation grounding, no public training data is available, and collecting such data would be particularly challenging, especially for multi-object grounding, due to the need for diversity. To address this, recent works have explored several reward-based fine-tuning approaches [17, 68, 7, 52, 70], but these methods require extensive computational resources. Instead of fine-tuning, we propose a test-time reward-guided framework that leverages a discriminative foundational model for guidance.
@@ 3 Origen
We present Origen, a zero-shot method for 3D orientation grounding in text-to-image generation. To the best of our knowledge, this is the first 3D orientation grounding method for multiple objects across diverse categories.
3.1 Problem Formulation and Overview
@@
As there is no training dataset consisting of diverse real-world images with accurate per-object orientation annotations, we propose a zero-shot approach that leverages test-time guidance from the recently proposed foundational model for image-based orientation estimation: OrientAnything [67]. In particular, given this orientation estimator $\mathcal{D}$ and a reward function $\mathcal{R}$ defined based on it (which we will discuss in detail in Sec. 3.2), our goal is to find a latent sample $\mathbf{x}$ that maximizes the reward of its corresponding image, expressed as $\mathcal{R}(\mathcal{F}_{\theta}(\mathbf{x},c),\mathcal{W},\Phi)$. For simplicity, we define the pullback of the reward function as ^$\hat{\mathcal{R}}=\mathcal{R}\circ\mathcal{F}_{\theta}$ and, unless otherwise specified, omit $\mathcal{W}$ and $\Phi$ in the notation, writing the objective simply as ^$\hat{\mathcal{R}}(\mathbf{x})$ in the following sections.
A straightforward approach to maximizing $\hat{\mathcal{R}}$ is gradient ascent, where the latent sample $\mathbf{x}$ is iteratively optimized. The update rule at each optimization step can be written as:
@@
where $\eta$ is a step size. However, this gradient ascent in the latent space may pose several challenges: (1) the latent sample $\mathbf{x}$ may get stuck in local maxima [69, 54] before achieving the desired orientation alignment, (2) the mode-seeking nature of gradient ascent can reduce sample diversity [26], and (3) $\mathbf{x}$ may deviate from the prior latent distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$, resulting in unrealistic images [17, 63]. Although the recent reward-guided noise optimization method (ReNO [16]) employs norm-based regularization [5, 56] to enforce the latent to be close to the prior distribution, it still suffers from local optima, leading to suboptimal orientation grounding results (see Sec. 4 for experimental comparisons and Appendix A for detailed analysis).
To address these issues, we reformulate the problem as a sampling problem rather than an optimization problem. Specifically, we aim to sample $\mathbf{x}$ from a target distribution $q^{*}$ that maximizes the expected reward, while ensuring $q$ remains close to the original latent distribution:
@@
where $\alpha\in\mathbb{R}$ is a constant that controls the regularization strength, and $D_{\text{KL}}(\,\cdot\,\|\,\cdot\,)$ is the Kullback-Leibler divergence [32].
This objective is closely related to those considered in existing fine-tuning-based approaches for reward maximization [53, 17, 35]. However, the key difference is that, while these methods define the target distribution $q^{*}$ for the output images, we define it for the latent samples, setting $q$ as the prior distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$. This formulation enables the derivation of an effective and simple sampling approach based on Langevin dynamics, which we discuss in Sec. 3.3.
We first introduce our reward function designed for 3D orientation grounding (Sec. 3.2) and propose reward-guided Langevin dynamics to effectively sample from our target distribution $q^{*}$ (Sec. 3.3). Lastly, we introduce reward-adaptive time rescaling to speed up sampling convergence by incorporating time rescaling (Sec. 3.4).
3.2 Orientation Grounding Reward
To define our reward function $\mathcal{R}$, we leverage a foundational orientation estimation model, OrientAnything [67], (henceforth denoted as $\mathcal{D}$) and measure how well the orientations of the objects estimated by $\mathcal{D}$ from the generated image align with the grounding orientation conditions $\Phi$. Since $\mathcal{D}$ represents the orientation as a probability distribution discretized into one-degree intervals, we define $\mathcal{R}$ based on the negative KL divergence between the estimated and target probability distributions of discretized orientations:
$\mathcal{R}(\mathbf{I},\mathcal{W},\Phi)=-\frac{1}{N}\sum_{i=1}^{N}\sum_{j\in\mathcal{S}}D_{\text{KL}}\Big{(}\mathcal{D}^{j}\big{(}\mathrm{Crop}(\mathbf{I},w_{i})\big{)}\,\Big{\|}\,\Pi(\phi_{i}^{j})\Big{)},$ (3)
[…]
@@ 3.3 Reward-Guided Langevin Dynamics
$d\mathbf{x}_{t}=\left(\frac{-\mathbf{x}_{t}+\frac{1}{\alpha}\nabla\hat{\mathcal{R}}(\mathbf{x}_{t})}{2}\right)dt+d\mathbf{w}_{t},$ (4)
@@
The proof of Proposition 1 is provided in Appendix B. This demonstrates that simulating the Langevin stochastic differential equation (SDE) (Eq. 4) in the latent space ensures samples $\mathbf{x}$ are drawn from the target distribution $q^{*}$, balancing reward maximization with proximity to the prior distribution (Eq. 2). Using Euler–Maruyama discretization [18], we express its discrete-time approximation as:
@@
$\mathbf{x}_{i+1}=\sqrt{1-\gamma}\,(\mathbf{x}_{i}+\gamma\eta\nabla\hat{\mathcal{R}}(\mathbf{x}_{i}))+\sqrt{\gamma}\epsilon_{i}.$ (6)
@@ 3.4 Reward-Adaptive Time Rescaling
$d\tilde{\mathbf{x}}_{\tau}$ $=\mathcal{G}(\hat{\mathcal{R}}(\tilde{\mathbf{x}}_{\tau}))\left(-\frac{1}{2}\tilde{\mathbf{x}}_{\tau}+\eta\nabla\hat{\mathcal{R}}(\tilde{\mathbf{x}}_{\tau})\right)d\tau$
$\quad+\frac{1}{2}\nabla\mathcal{G}(\hat{\mathcal{R}}(\tilde{\mathbf{x}}_{\tau}))d\tau+\sqrt{\mathcal{G}(\hat{\mathcal{R}}(\tilde{\mathbf{x}}_{\tau}))}d\tilde{\mathbf{w}}_{\tau},$ (7)
@@
$\mathbf{x}_{i+1}=\;$ $\sqrt{1-\gamma(\mathbf{x}_{i})}\Bigl{(}\mathbf{x}_{i}+\gamma(\mathbf{x}_{i})\eta\,\nabla\hat{\mathcal{R}}(\mathbf{x}_{i})\Bigr{)}$
$+\frac{1}{2}\gamma(\mathbf{x}_{i})\nabla\log\mathcal{G}(\hat{\mathcal{R}}(\mathbf{x}_{i}))+\sqrt{\gamma(\mathbf{x}_{i})}\,\epsilon_{i}.$ (8)
Regarding the design of the monitor function, existing work [38] suggests setting the step size inversely proportional to the squared norm of the drift coefficients in Eq. 4. However, this approach requires computing the Hessian of the reward function when evaluating the correction term, which is computationally expensive. Alternatively, we propose the following simple, reward-adaptive monitor function:
@@
where the hyperparmeters $s_{\text{min}},\,s_{\text{max}},$ and $k$ are set to $\frac{1}{3}$, $\frac{4}{3}$, and $\frac{6}{5}$ in our experiments, respectively. This function adaptively scales the step size by assigning smaller steps in high-reward and larger steps in low-reward, thereby improving convergence speed and accuracy. Please refer to Fig. B in Appendix that provides the visualization of our monitor function $\mathcal{G}(\hat{\mathcal{R}}(\mathbf{x}))$.
In Fig. 2, we present a toy experiment demonstrating the effectiveness of our Langevin dynamics and reward-adaptive time rescaling compared to gradient ascent with regularization (ReNO [16]). See Appendix D for details on the toy experiment setup. The top row shows the ground truth target latent distribution $q^{*}$ (leftmost) alongside latent samples generated by different methods, and the bottom row displays the corresponding data distributions. While ReNO [16] fails to accurately capture the target distribution due to its mode-seeking behavior (2nd column, as discussed in Sec. 3.1), our method successfully aligns with it (3rd column), and time rescaling further accelerates convergence (4th column).
Our optimization procedure is outlined in Alg. 1. Note that setting $\mathcal{G}(\hat{\mathcal{R}}(\mathbf{x}))=1$ reverts the method to the uniformly time-rescaled form.
@@
1:$c$ (Prompt), $\mathcal{W}$ (Object phrase set), $\Phi$ (Grounding orientation set), $\mathcal{F}_{\theta}$ (One-step T2I model), $\mathcal{R}$ (Reward function), $\mathcal{G}$ (Monitor function), $M$ (# of optimization steps), $\eta$ (Gradient scaling factor), $\gamma$ (Step size)
2:$\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
@@
4:     ^$\hat{\mathcal{R}}_{i}=\mathcal{R}(\mathcal{F}_{\theta}(\mathbf{x}_{i},c),\mathcal{W},\Phi)$triangleright Reward computation
@@
7:     $\mathbf{x}_{i+1}=\sqrt{1-\gamma_{i}}\Bigl{(}\mathbf{x}_{i}+\gamma_{i}\eta\nabla\hat{\mathcal{R}}_{i}\Bigr{)}$
8:               $+\frac{1}{2}\gamma_{i}\nabla\log\mathcal{G}(\hat{\mathcal{R}}_{i})$ + $\sqrt{\gamma_{i}}\epsilon_{i}$triangleright Update step
9:end for
10:return $\mathcal{F}_{\theta}(\mathbf{x}_{M},c)$
@@
Figure 2: Toy experiment results. The top row shows latent space samples (red), while the bottom row shows the corresponding data space samples (blue). From left to right, each column represents: (1) the ground truth target distribution from Eq. (2); (2) results of ReNO [16]; (3) results of ours with uniform time scaling; and (4) results of ours with reward-adaptive time rescaling.

### v2

@@ Abstract
Figure 1: 3D orientation-grounded text-to-image generation results of Origen. We present Origen, the first zero-shot method for 3D orientation grounding in text-to-image generation across multiple objects and diverse categories. Origen generates high-quality images that are accurately aligned with the grounding orientation conditions (colored arrows) and the input text prompts.
@@ 1 Introduction
Since no existing method has quantitatively evaluated 3D orientation grounding in text-to-image generation (except for user studies by [18]), we curate a benchmark based on the MS-COCO dataset [23], mixing and matching object classes and orientations to create images with single or multiple orientation-grounded objects. We demonstrate that Origen significantly outperforms previous orientation-conditioned image generative models [18, 17] on both our benchmark and user studies. Since prior models cannot condition on multiple objects (whereas Origen can, as shown in Fig. 1), comparisons are conducted under single-object conditioning. Additionally, we perform experiments to further validate the superior performance of our method over text-to-image generative models with orientation-specific prompts and other training-free guided sampling strategies.
@@
  • We also propose a reward-adaptive time rescaling method that accelerates convergence.
@@ 2 Related Work
A number of training-free methods have been proposed for guided generation tasks. DPS [32], MPGD [33], and Pi-GDM [34] update the noisy data point at each step using a given reward function. FreeDoM [35] takes this further by introducing rewinding, where intermediate data points are regenerated by reversing the generative denoising process. The core principle of this approach is leveraging the posterior mean from a noisy image at an intermediate step via Tweedie’s formula [36]. The expected future reward can also be computed from the posterior mean, allowing gradient ascent to update the noisy image. However, a key limitation of these methods is that the posterior mean from a diffusion model is often too blurry to accurately predict future rewards. While distilled or fine-tuned diffusion [37, 38, 39, 40] and flow models [41, 42] can mitigate this issue, their straightened trajectories lead to insufficiently small updates to the image during gradient ascent at intermediate timesteps.
Recent approaches [22, 43, 44, 45] attempt to address this limitation by updating the initial noise rather than the intermediate noise. Notably, ReNO [22] is an optimization method based on a one-step generative model to efficiently iterate the initial noise update through one-step generation and future reward computation. However, gradient ascent with respect to the initial noise often suffers from local optima and leads to deviations from real images, even with heuristic regularization [46]. To overcome this limitation, we propose a novel sampling-based approach rather than an optimization-based one, leveraging Langevin dynamics to effectively balance reward maximization and realism.
@@
For controlling image generation, ControlNet [28] and IP-Adapter [47] are commonly used to utilize a pretrained generative model to control for various conditions, though they require training data. For 3D orientation grounding, no public real-world training data is available, and collecting such data would be particularly challenging, especially for multi-object grounding, due to the need for diversity. To address this, recent works have introduced RL-based reward fine-tuning approaches [48, 49, 50, 51, 52]. However, these methods require substantial computational resources and, more importantly, cause significant deviation from the pretrained data distribution when trained with task-specific rewards, leading to degraded image quality and reduced diversity [53]. Hence, instead of fine-tuning, we propose a test-time reward-guided framework that leverages a discriminative foundational model for guidance.
@@ 3 Origen
We present Origen, a zero-shot method for 3D orientation grounding in text-to-image generation. To the best of our knowledge, this is the first 3D orientation grounding method for multiple objects across open-vocabulary categories.
3.1 Problem Definition and Overview
@@
Why Test-Time Guidance?
Supervised methods are not applicable to this task due to the lack of training datasets containing diverse real-world images with per-object orientation annotations. Previous supervised approaches [18, 17] have therefore been limited to single-object scenarios within a narrow set of categories. To address this, we propose a training-free approach that leverages recent foundational models to design a reward function: GroundingDINO [55] for object detection and OrientAnything [20] for orientation estimation. While this reward function could also be used in recent RL-based self-supervised methods for fine-tuning the image generative model, such techniques not only demand substantial computational resources but also lead to degraded image quality when trained with such a task-specific reward [53]. To this end, we propose a novel training-free approach that effectively avoids image quality degradation while maximizing orientation reward at test time.
In particular, given a reward function $\mathcal{R}$ defined based on GroundingDINO and OrientAnything (which we will discuss in detail in Sec. 3.2), our goal is to find a latent sample $\mathbf{x}$ that maximizes the reward of its corresponding image, expressed as $\mathcal{R}(\mathcal{F}_{\theta}(\mathbf{x},c))$. For simplicity, we define the pullback of the reward function as ^$\hat{\mathcal{R}}=\mathcal{R}\circ\mathcal{F}_{\theta}$ and use this notation throughout.
Why Based on a One-Step Generative Model?
Despite extensive research on test-time reward-guided generation using pretrained diffusion models [32, 33, 35], we find that the main challenge stems from the multi-step nature of these models. Reward-guided generation requires computing rewards for the expected final output at intermediate denoising steps and applying gradient ascent. However, in early stages, the expected output is too blurry to provide meaningful guidance, while in later stages—when the output becomes clearer—gradient updates have minimal effect. In contrast, a one-step model generates a clear image directly from a prior sample, enabling more effective guidance. Comparative results are presented in Section 4.
Our Technical Contribution.
The gradient ascent approach for maximizing the future reward $\hat{\mathcal{R}}$ can also be applied to the one-step model, using the following update rule on the latent $\mathbf{x}_{i}$ at each iteration $i$:
@@
where $\eta$ is a step size. However, this gradient ascent in the latent space pose several challenges: (1) the latent sample $\mathbf{x}$ may get stuck in local maxima [56, 57] before achieving the desired orientation alignment, (2) the mode-seeking nature of gradient ascent can reduce sample diversity [58], and (3) $\mathbf{x}$ may deviate from the prior latent distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$, resulting in unrealistic images [48, 59]. Although the recent reward-guided noise optimization method (ReNO [22]) employs norm-based regularization [43, 60] to enforce the latent to be close to the prior distribution, it still suffers from local optima, leading to suboptimal orientation grounding results (see comparisons in Sec. 4 and further analysis in Appendix A).
Our key idea to address this is to reformulate the problem as a sampling problem rather than an optimization problem. Specifically, we aim to sample $\mathbf{x}$ from a target distribution $q^{*}$ that maximizes the expected reward, while ensuring $q$ remains close to the original latent distribution:
@@
where $\alpha\in\mathbb{R}$ is a constant that controls the regularization strength, and $D_{\text{KL}}(\,\cdot\,\|\,\cdot\,)$ is the Kullback-Leibler divergence [61]. This objective is closely related to those used in fine-tuning-based approaches for reward maximization [62, 48, 63]. However, the key difference is that, while these methods define the target distribution $q^{*}$ for the output images, we define it for the latent samples, setting $q$ as the prior distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$. This formulation leads to our key contribution: a novel, simple, and effective sampling method based on Langevin dynamics, discussed in Sec. 3.3.
Below, we first introduce our reward function designed for 3D orientation grounding (Sec. 3.2) and propose reward-guided Langevin dynamics to effectively sample from our target distribution $q^{*}$ (Sec. 3.3). Lastly, we introduce reward-adaptive time rescaling to speed up sampling convergence by incorporating time scheduling (Sec. 3.4).
3.2 Multi-Object Orientation Grounding Reward
To define our reward function $\mathcal{R}$, we first detect the described objects in the generated image using GroundingDINO [55] with the provided object phrases $\mathcal{W}$, yielding cropped object regions $\mathrm{Crop}(\mathbf{I},w_{i})$ for each phrase $w_{i}$. We then use OrientAnything [20] to assess how well the orientations of the detected objects align with the specified grounding conditions $\Phi$. OrientAnything represents object orientations as categorical probability distributions over one-degree bins for each Euler angle. Let $\mathcal{D}_{j}(\cdot)$ denote the predicted distribution for the $j$-th angle. Following the convention from OrientAnything, we define the target distribution $\Pi(\phi_{i,j})$ as a discretized Gaussian centered at the reference angle $\phi_{i,j}$. The reward function $\mathcal{R}$ is then computed as the negative KL divergence between the predicted and target distributions, summed across all angles and all objects:
$\mathcal{R}(\mathbf{I},\mathcal{W},\Phi)=-\frac{1}{N}\sum_{i=1}^{N}\sum_{j\in\mathcal{S}}D_{\text{KL}}\Big{(}\mathcal{D}_{j}\big{(}\mathrm{Crop}(\mathbf{I},w_{i})\big{)}\,\Big{\|}\,\Pi(\phi_{i,j})\Big{)}.$ (3)
@@ 3.3 Reward-Guided Langevin Dynamics
$d\mathbf{x}_{t}=\left(\frac{-\mathbf{x}_{t}+\frac{1}{\alpha}\nabla\hat{\mathcal{R}}(\mathbf{x}_{t})}{2}\right)dt+d\mathbf{w}_{t}$ (4)
@@
Proof.
See Appendix B. ∎
This demonstrates that simulating the Langevin stochastic differential equation (SDE) (Eq. 4) in the latent space ensures samples $\mathbf{x}$ are drawn from the target distribution $q^{*}$, balancing reward maximization with proximity to the prior distribution (Eq. 2). Using Euler–Maruyama discretization [64], we express its discrete-time approximation as:
@@
$\mathbf{x}_{i+1}=\sqrt{1-\gamma}\,\mathbf{x}_{i}+\gamma\eta\nabla\hat{\mathcal{R}}(\mathbf{x}_{i})+\sqrt{\gamma}\epsilon_{i}.$ (6)
@@ 3.4 Reward-Adaptive Time Rescaling
$d\tilde{\mathbf{x}}_{\tau}$ $=\mathcal{G}(\hat{\mathcal{R}}(\tilde{\mathbf{x}}_{\tau}))\left(-\frac{1}{2}\tilde{\mathbf{x}}_{\tau}+\eta\nabla\hat{\mathcal{R}}(\tilde{\mathbf{x}}_{\tau})\right)d\tau+\frac{1}{2}\nabla\mathcal{G}(\hat{\mathcal{R}}(\tilde{\mathbf{x}}_{\tau}))d\tau+\sqrt{\mathcal{G}(\hat{\mathcal{R}}(\tilde{\mathbf{x}}_{\tau}))}d\tilde{\mathbf{w}}_{\tau},$ (7)
@@
$\mathbf{x}_{i+1}=\;$ $\sqrt{1-\gamma(\mathbf{x}_{i})}\mathbf{x}_{i}+\gamma(\mathbf{x}_{i})\eta\,\nabla\hat{\mathcal{R}}(\mathbf{x}_{i})+\frac{1}{2}\gamma(\mathbf{x}_{i})\nabla\log\mathcal{G}(\hat{\mathcal{R}}(\mathbf{x}_{i}))+\sqrt{\gamma(\mathbf{x}_{i})}\,\epsilon_{i}.$ (8)
Regarding the design of the monitor function, existing work [65] suggests setting the step size inversely proportional to the squared norm of the drift coefficients in Eq. 4. However, this approach requires computing the Hessian of the reward function for the correction term, which is computationally heavy. Alternatively, we propose following simple, reward-adaptive monitor function:
@@
where the hyperparmeters $s_{\text{min}},\,s_{\text{max}},$ and $k$ are set to $\frac{1}{3}$, $\frac{4}{3}$, and $0.3$ in our experiments, respectively. This function adaptively scales the step size by assigning smaller steps in high-reward and larger steps in low-reward, thereby improving convergence speed and accuracy. Please refer to Fig. 5 in Appendix that provides the visualization of our monitor function $\mathcal{G}(\hat{\mathcal{R}}(\mathbf{x}))$.
In Fig. 2, we present a toy experiment demonstrating the effectiveness of our Langevin dynamics and reward-adaptive time rescaling compared to gradient ascent with regularization (ReNO [22]). See Appendix D for details on the toy experiment setup. The top row shows the ground truth target latent distribution $q^{*}$ (leftmost) alongside latent samples generated by different methods, and the bottom row displays the corresponding data distributions. While ReNO [22] fails to accurately capture the target distribution due to its mode-seeking behavior (2nd column, as discussed in Sec. 3.1), our method successfully aligns with it (3rd column), and time rescaling further accelerates convergence (4th column). Our sampling procedure is outlined in Alg. 1. Note that setting $\mathcal{G}(\hat{\mathcal{R}}(\mathbf{x}))=1$ reverts the method to the uniformly time-rescaled form.
@@
1:$c$ (Prompt), $\mathcal{W}$ (Object phrase set), $\Phi$ (Grounding orientation set), $\mathcal{F}_{\theta}$ (One-step T2I model), $\mathcal{R}$ (Reward), $\mathcal{G}$ (Monitor), $M$ (# Steps), $\eta$ (Scale), $\gamma$ (Step size)
2:$\text{Initialize }\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),\mathcal{R}_{*}=-\inf$
@@
4:    ^$\hat{\mathcal{R}}_{i}=\mathcal{R}(\mathcal{F}_{\theta}(\mathbf{x}_{i},c),\mathcal{W},\Phi)$triangleright Reward Calc.
@@
7:    $\mathbf{x}_{i+1}=\sqrt{1-\gamma_{i}}\mathbf{x}_{i}+\gamma_{i}\eta\nabla\hat{\mathcal{R}}_{i}$$+\frac{1}{2}\gamma_{i}\nabla\log\mathcal{G}(\hat{\mathcal{R}}_{i})$
8:               + $\sqrt{\gamma_{i}}\epsilon_{i}$triangleright Update step
9:    if ^$\hat{\mathcal{R}}_{i}>\mathcal{R}_{*}$ then
10:        =^$\mathcal{R}_{*}=\hat{\mathcal{R}}_{i},\mathbf{x}_{*}=\mathbf{x}_{i}$
11:    end if
12:end for
13:return $\mathcal{F}_{\theta}(\mathbf{x}_{*},c)$
@@
Figure 2: Toy experiment results. Top: latent space samples (blue); bottom: data space samples (red). Gray dots show the original distribution without reward guidance. From left to right: (1) ground truth target distribution from Eq. 2, (2) results of ReNO [22], (3) results of ours with uniform time scaling, and (4) results of ours with reward-adaptive time rescaling.